OpenStudy (anonymous):

divide & simplify: sqrt20 / sqrt5

jimthompson5910 (jim_thompson5910):

sqrt(20)/sqrt(50=sqrt(20/5)=sqrt(4)=2 So \[\large \frac{\sqrt{20}}{\sqrt{5}}=2\]

jimthompson5910 (jim_thompson5910):

oops made a typo at top of the line, should be sqrt(20)/sqrt(5)

jimthompson5910 (jim_thompson5910):

answer is still the same though